# Multi-Agent Deep Reinforcement Learning for Optimized Multi-UAV Coverage and Power-Efficient UE Connectivity

### v1

@@ I Introduction
Compared to centralized DQN and equal power allocation, the proposed decentralized MADDPG strategy improves UE service efficiency by up to 2.07 times and 8.84 times, respectively.
@@ II Related Work
The optimization of power allocation and deployment strategies in UAV-assisted systems has garnered significant attention, particularly in mobile edge computing (MEC) scenarios and energy-efficient UAV operations. Several studies have employed RL techniques to address these challenges, focusing on trajectory design, task offloading, and energy management. [1] proposed various MEC frameworks utilizing MADDPG algorithms to optimize task scheduling, trajectory planning, and resource allocation. These studies demonstrated significant improvements in energy efficiency, reduced task processing delays, and fairness in resource distribution across UAVs. Additionally, [14] extended these approaches to ensure geographical and load fairness while optimizing energy consumption for UAV-assisted MEC networks. For energy-efficient UAV path planning, [15, 5] introduced MADDPG-based algorithms that minimized energy usage through techniques like pruning and optimization of neuron layers, as well as by addressing eavesdropping threats in MEC systems with ground-based jamming. Similarly, [16] developed a decentralized trajectory and power control algorithm for UAV base stations (ABSs), achieving high UE satisfaction and operational scalability using centralized training with decentralized execution paradigms. In air-ground collaborative networks, [17] proposed multi-UAV systems leveraging Lyapunov optimization and MADDPG for adaptive task offloading, service instance management, and resource allocation. These approaches minimized energy consumption and economic expenditure, demonstrating fast convergence and superior cost efficiency compared to baseline methods. Dynamic and adaptive UAV operations were further explored in [18, 4], where advanced algorithms like MADDPG-LC, Multi-Agent Proximal Policy Optimization (MAPPO), and PPO2-based DRL were employed for dynamic trajectory control, cooperative UAV swarm management, and 3-D trajectory design. These studies highlighted improved energy efficiency, faster convergence, and robustness in addressing flight dynamics and disaster recovery scenarios. [6] employs DQN and DDPG to address bandwidth and power allocation for a single UAV operating in a static, interference-free environment. However, cooperation among multiple UAVs is essential for more complex scenarios.
@@
In contrast to previous works that primarily focus on energy efficiency, task offloading, and trajectory planning using RL techniques such as MADDPG, MAPPO, and DDPG, our contributions differentiate our work from the related literature by addressing the challenges of optimal spatial UAV deployment and dynamic interference management by optimally allocating power to UEs in multi-UAV-assisted networks. While the existing literature provides a robust foundation for UAV-assisted wireless communication, several key areas remain for further investigation, such as extending current approaches to address dense urban environments and large-scale UAV deployments and enhanced coverage efficiency by optimizing power usage without compromising network performance.
@@ III-A System Model
UAVs enhance wireless coverage by providing flexible deployment and connectivity in challenging environments. We consider a multi-UAV-assisted communication system, where $N$ UEs are uniformly distributed across a two-dimensional square field $\Psi$ with sides of $L$ meters. The square field consists of $100\times 100$ grids with each grid cell being a square of side $l=L/100$ meters. In this setup, illustrated in Fig. 1, these users are grouped in different clusters and communicate with a dedicated UAV for each cluster. The number of UAVs deployed are equal to the number of UEs’ clusters. Note that the user clustering and UAVs deployment approach is discussed in the next section. Each UE $UE_{i}$ is identified by $i\in I{\triangleq}{1,2,\ldots,N}$, and their respective positions are defined by coordinates $(x_{i},y_{i},0)$ relative to the left lower vertex of $\Psi$, $(0,0,0)$. The location of each UAV $UAV_{j}$ is represented by coordinates $(x_{j},y_{j},h)$ in three-dimensional space, where $h$ represents the hovering height of all UAVs.
Consider a designated UE $UE_{i}$ depicted in Fig. 1, situated at a horizontal distance $d_{i.j}\triangleq\sqrt{(x_{j}-x_{i})^{2}+(y_{j}-y_{i})^{2}}$ from the associated UAV $UAV_{j}$, and the elevation angle of the UAV $UAV_{j}$ to that user is $\theta_{i,j}$ rad. For simplicity, we utilize Euclidean distance metrics in our analysis. Given that UAVs maintain an altitude of $h$ meters above the field $\Psi$, the distance between UE $UE_{i}$ and the UAV $UAV_{j}$ can be calculated as $r_{i,j}\triangleq\sqrt{d_{i,j}^{2}+h^{2}}=\frac{h}{sin(\theta_{i})}$. All the UAVs and users are assumed to be equipped with a single antenna.
@@ V-A Environment and MADDPG Training Parameters
Unless otherwise stated all the environmental parameters used in the simulation setup are provided in Table I. The hyperparameters in MADDPG training are selected to ensure stable and efficient learning. A replay buffer of 100,000 experiences supports off-policy learning, while a batch size of 64 stabilizes updates. A learning rate of 0.0001 is chosen to prevent drastic weight updates and enhance convergence stability, and a discount factor $\gamma=0.95$ balances short and long-term rewards. A soft update rate $\tau=0.01$ ensures smooth policy updates, while an exploration noise of $\sigma_{\text{noise}}=0.2$ promotes exploration, preventing premature convergence to suboptimal policies. The selected values of all hyperparameters are summarized in Table II. The simulations have been conducted over 10 different seeds, and all figures present the average results across these 10 runs, with $95\%$ confidence interval (CI) bars indicating variability.
@@
Figure 4: Num. of Served UEs ($N=60$ UEs, $R_{th}=30$ Mbps)
Figure 5: MADDPG performance for different cell scales ($N=30$ UEs, $R_{th}=30$ Mbps)
@@ V-B1 Training Convergence
Fig. 2 shows the convergence behavior of MADDPG training over $500$ time-steps for $3$, $5$, and $7$ clusters. The reward starts low and steadily increases, stabilizing after $100$ time-steps. Minor drops likely result from policy updates or exploration-exploitation trade-offs. Training with 7 clusters yields the most stable policy with minimal variance and the highest reward, suggesting that more clusters enhance learning through richer interactions and exploration. Additionally, a higher number of clusters provides more UAV resources to serve extra users, leading to improved rewards compared to fewer clusters.

### v2

@@ I Introduction
Compared to centralized DQN and equal power allocation, the proposed decentralized MADDPG strategy improves UE coverage efficiency by serving up to 2.07 times and 8.84 times more users, respectively.
@@ II Related Work
The optimization of power allocation and deployment strategies in UAV-assisted systems has garnered significant attention, particularly in mobile edge computing (MEC) scenarios and energy-efficient UAV operations. Several studies have employed RL techniques to address these challenges, focusing on trajectory design, task offloading, and energy management. [1] proposed various MEC frameworks utilizing MADDPG algorithms to optimize task scheduling, trajectory planning, and resource allocation. These studies demonstrated significant improvements in energy efficiency, reduced task processing delays, and fairness in resource distribution across UAVs. Additionally, [14] extended these approaches to ensure geographical and load fairness while optimizing energy consumption for UAV-assisted MEC networks. For energy-efficient UAV path planning, [15, 5] introduced MADDPG-based algorithms that minimized energy usage through techniques like pruning and optimization of neuron layers, as well as by addressing eavesdropping threats in MEC systems with ground-based jamming. The study presented in [16] focuses on the joint optimization of content caching probability, resource allocation, and UAV flight trajectory. To achieve this, the authors propose a MADDPG-based Resource allocation and UAV trajectory Optimization (MRFO) algorithm to maximize the overall system energy efficiency. In air-ground collaborative networks, [17] proposed multi-UAV systems leveraging Lyapunov optimization and MADDPG for adaptive task offloading, service instance management, and resource allocation. These approaches minimized energy consumption and economic expenditure, demonstrating fast convergence and superior cost efficiency compared to baseline methods. Dynamic and adaptive UAV operations were further explored in [18, 4], where advanced algorithms like MADDPG-LC, Multi-Agent Proximal Policy Optimization (MAPPO), and PPO2-based DRL were employed for dynamic trajectory control, cooperative UAV swarm management, and 3-D trajectory design. These studies highlighted improved energy efficiency, faster convergence, and robustness in addressing flight dynamics and disaster recovery scenarios. [6] employs DQN and DDPG to address bandwidth and power allocation for a single UAV operating in a static, interference-free environment. However, cooperation among multiple UAVs is essential for more complex scenarios.
@@
In contrast to previous works that primarily focus on energy efficiency, task offloading, and trajectory planning using RL techniques such as MADDPG, MAPPO, and DDPG, our contributions differentiate our work from the related literature by addressing the challenges of optimal spatial UAV deployment and dynamic interference management by optimally allocating power to UEs in multi-UAV-assisted networks to enhance coverage efficiency by maximizing the number of users served, rather than prioritizing energy efficiency. While the existing literature provides a robust foundation for UAV-assisted wireless communication, several key areas remain for further investigation, such as extending current approaches to address dense urban environments and large-scale UAV deployments and enhanced coverage efficiency by optimizing power usage without compromising network performance.
@@ III-A System Model
UAVs enhance wireless coverage by providing flexible deployment and connectivity in challenging environments. We consider a multi-UAV-assisted communication system, where $N$ UEs are uniformly distributed across a two-dimensional square field $\Psi$ with sides of $L$ meters. The square field consists of $100\times 100$ grids with each grid cell being a square of side $l=L/100$ meters. In this setup, illustrated in Fig. 1, these users are grouped in different clusters and communicate with a dedicated UAV for each cluster. The number of UAVs deployed are equal to the number of UEs’ clusters. Note that the user clustering and UAVs deployment approach is discussed in the next section. Each UE $UE_{i}$ is identified by $i\in I{\triangleq}{1,2,\ldots,N}$, and their respective positions are defined by coordinates $(x_{m},y_{,}0)$ relative to the left lower vertex of $\Psi$, $(0,0,0)$. The location of each UAV $UAV_{j}$ is represented by coordinates $(x_{q},y_{q},h_{q})$ in three-dimensional space, where $h$ represents the hovering height of all UAVs.
Consider a designated UE $i$ depicted in Fig. 1, situated at a horizontal distance $d_{i.j}\triangleq\sqrt{(x_{j}-x_{i})^{2}+(y_{j}-y_{i})^{2}}$ from the associated UAV $UAV_{j}$, and the elevation angle of the UAV $UAV_{j}$ to that user is $\theta_{i,j}$ rad. For simplicity, we utilize Euclidean distance metrics in our analysis. Given that UAVs maintain an altitude of $h$ meters above the field $\Psi$, the distance between UE $UE_{i}$ and the UAV $UAV_{j}$ can be calculated as $r_{i,j}\triangleq\sqrt{d_{i,j}^{2}+h^{2}}=\frac{h}{sin(\theta_{i})}$. All the UAVs and users are assumed to be equipped with a single antenna.
@@ V-A Environment and MADDPG Training Parameters
Unless otherwise stated all the environmental parameters used in the simulation setup are provided in Table I. The hyperparameters in MADDPG training are selected to ensure stable and efficient learning. A replay buffer of 100,000 experiences supports off-policy learning, while a batch size of 64 stabilizes updates. A learning rate of 0.0001 is chosen to prevent drastic weight updates and enhance convergence stability, and a discount factor $\gamma=0.95$ balances short and long-term rewards. A soft update rate $\tau=0.01$ ensures smooth policy updates, while an exploration noise of $\sigma_{\text{noise}}=0.2$ promotes exploration, preventing premature convergence to suboptimal policies. The selected values of all hyper-parameters are summarized in Table II. In our setup, training typically requires over 500 episodes where each episode includes up to 500 time steps, resulting in several hours of GPU computation time. The simulations have been conducted over 10 different seeds, and all figures present the average results across these 10 runs, with $95\%$ confidence interval (CI) bars indicating variability.
@@
Figure 4: Performance comaprison for $N=60$ UEs ($R_{th}=30$ Mbps)
Figure 5: MADDPG performance for different cell scales (N=30, $R_{th}$=30 Mbps)
@@ V-B1 Training Convergence
Fig. 2 shows the convergence behavior of MADDPG training over $500$ time-steps for $3$, $5$, and $7$ clusters. The reward starts low and steadily increases, stabilizing after $100$ time-steps. Minor drops likely result from policy updates or exploration-exploitation trade-offs. Training with 7 clusters yields the most stable policy with minimal variance and the highest reward, suggesting that more clusters enhance learning through richer interactions and exploration.